# Probing supermassive black hole scalarization with Pulsar Timing Arrays

### v1

@@ Abstract
Scalar-tensor theories with a scalar field coupled to the Gauss-Bonnet invariant can evade no-hair theorems and allow for non-trivial scalar profiles around black holes. This coupling is characterized by a length scale $\lambda$, which, in an effective field theory perspective, sets the threshold below which deviations from General Relativity become significant. LIGO/VIRGO constraints indicate $\lambda$ is small, implying supermassive black holes should not scalarize. However, recent work suggests that scalarization can occur within a narrow window of masses, allowing supermassive black holes to scalarize, while leaving LIGO/VIRGO sources unaffected. We explore the impact of this scenario on the stochastic gravitational wave background recently observed by Pulsar Timing Arrays. We find that scalarization can alter the characteristic strain produced by circularly inspiralling SMBH binaries and that current data shows a marginal preference for a non-zero $\lambda$. However, similar signatures could arise from astrophysical effects such as orbital eccentricity or environmental interactions, emphasizing the need for improved modeling and longer observations to discriminate among the different scenarios.
@@ I Introduction
On the other hand, it has been shown that most of the interaction terms admitted in (ghost-free) ST gravity do not affect black hole (BH) solutions at the background level. In other words, BHs in most ST theories are exactly the same as in GR, meaning that they are fully determined by mass, spin and, in principle, electric charge, while the additional scalar DOF provides no further “charge” to the BH. This fact is related to the well-known no-hair conjecture and has been rigorously formalized in some specific cases through no-hair theorems [9, 10, 11, 12, 13]. Nevertheless, such theorems rely on a set of crucial assumptions that cannot always be taken for granted. In particular, it has been shown that a specific interaction term, among all those leading to second-order equations of motion (EOM), can violate all known no-hair theorems, allowing for BH solutions with a non-trivial scalar profile [14, 15]. This term is given by $\lambda^{2}F(\varphi)\mathcal{G}$, where $\lambda$ is a coupling constant with dimension of a length, $\varphi$ is the scalar field, $F$ is an analytic function, and $\mathcal{G}$ is the Gauss-Bonnet (GB) invariant
@@
The interesting properties of this model include BH solutions coupled to the scalar field via the so-called scalar charge [14, 15, 16, 17, 18], which has been used to constrain the coupling parameter $\lambda$ using Gravitational Waves (GWs) observations from the LIGO/Virgo collaboration [2]. An interesting feature of this model is that the best constraints on $\lambda$ actually come from the smaller BHs from which we observe the GWs. This is tied to the dimensionful nature of the coupling: the smaller the BH scale, the better the constraint on $\lambda$. In fact, one can view interaction terms coming with a dimensionful coupling constant $\lambda$, such as the scalar-GB coupling, as irrelevant operators in an effective field theory (EFT) perspective [19]. This implies that the contribution of this operator is suppressed at scales that are much larger than $\lambda$, whereas deviations from GR are expected to become increasingly significant at smaller scales. For this reason, it is usually assumed that supermassive BHs (SMBHs) are the same as in GR, even in presence of the scalar-GB coupling.
@@
The EFT interpretation of ST theories can also be made more concrete, as it can be shown that some specific operators arise as the low-energy limit of quantum gravity theories. This is the case for some scalar-GB couplings, that emerge in the low-energy limit of string theories [20]. However, in principle, one can suppose that high-energy effects, at a given scale, lead BHs back to the GR configuration, providing a well-delimited range of scales in which deviations from GR can occur. This happens, for instance, in a recently proposed model [18], in which SMBHs can possess a scalar charge while lower mass BHs are the same as in GR.
@@
It is therefore reasonable to envision departures from GR in the signals that we receive from SMBHs, which, when coupled in binary systems, are considered the most likely source of the recently observed stochastic GW background (SGWB). Indeed, in recent years, Pulsar Timing Array (PTA) collaborations reported evidence for a temporally-correlated stochastic process leaving its imprint on the times of arrival (TOAs) of radio beams coming from a network of monitored millisecond pulsars in our Galaxy [21, 22, 23, 24, 25]. Furthermore, the correlations observed among different pulsars were found to be consistent with the Hellings-Downs (HD) pattern [26], as expected for a SGWB. Although the observed data can also be explained with Early Universe processes (see e.g. Refs. [27, 28]), the most likely interpretation consists in the incoherent superposition of GWs emitted by a population of inspiralling SMBH binaries (SMBHBs). Assuming circular orbits — which we restrict to in the following — the characteristic strain spectrum associated with GW-driven orbital decay is expected to follow a power-law scaling as $f^{-n}$, with $n=2/3$, where $f$ denotes the emitted GW frequency [29, 30]. Eccentric orbits as well as the presence of any additional non-GW energy loss channel - such as dynamical friction or, in our case, scalar radiation- would accordingly modify the spectral index $n$ [31, 32, 33, 34, 35, 36, 37].
On the other hand, the predicted angular correlation pattern is a property of the polarization of GWs at the Earth location, and will essentially coincide with the HD curve whenever - as we consider in the following - the scalar field is not coupled to matter.
@@
In this article, we hence study the modification of the SGWB predicted by an assembly of circular SMBHBs evolving under the ST theory of gravity studied in Ref. [18]. In Sec. II, we summarize the framework of black hole scalarization and lay the foundations for the analysis developed throughout the paper. In Sec. III, we review how to construct the SGWB from an astrophysical population of SMBHBs, and we compare the resulting characteristic strain in GR with that predicted by the ST theory described in Sec. II. In Sec. IV, we analyze the EPTA DR2New dataset, comprising the most recent 10.3 yr of data of the EPTA collaboration, and the NANOGrav 15yr data, looking for signatures of non trivial SMBHBs scalarization. Finally, in Sec. V, we summarize our findings and conclude with a brief discussion of their implications and possible directions for future work. Additional technical details and supplementary plots are provided in the Appendices. In the following, we will work in natural units $c=1$, while keeping expressed the gravitational coupling constant $G$, and using the mostly plus signature of the metric tensor (,+,+,+)(-,+,+,+)( - , + , + , + ).
@@ II.1 Motivation
Notice that the shift-symmetric models exhibit characteristic features, like finite-radius singularities, whose location depends on the coupling $\lambda$ [14, 15]. The presence of these singularities, if one requires them to be always located within the event-horizon, imposes a lower bound on the BH size, that translates into an observational constraint on the coupling $\lambda$ for this theory [42, 43]. However, it is worth mentioning that the shift-symmetric EsGB theory, when requiring that it produces observable deviations from GR at astrophysical scales, presents issue with causality [44]. In the dilatonic case, one also encounters a lower BH mass limit, set by the dimensionful coupling $\lambda$, so that the best experimental constraints can be put from low mass BH observations [45].
@@
with $\tilde{\lambda}$ being another dimensionful coupling constant. In this model, to ensure that the Schwarzschild BH still solve the EOMs, one has to restrict to the case $\varphi=0$. In the $\tilde{\lambda}=0$ case, the Schwarzschild solution is admitted with any constant $\varphi_{0}$, as $\mathcal{G}$ yields a trivial contribution to the EOM, being it a surface term. In the more general case $\tilde{\lambda}\neq 0$, the condition for the tachyonic instability to occurr is now given by
@@
This practically demonstrates that higher-order terms in EsGB can, in principle, protect BHs from scalarization at small scales, where the standard exponential model described in Ref. [16] is affected by hyperbolicity loss, while allowing for this phenomenon to occurr at larger scales. For this reason, the constraints on scalar charges provided by LIGO-Virgo do not necessarily imply that scalarization should be excluded for SMBHs.
In the following, we will be agnostic about the specific model, and just assume that the scalar field is coupled to some positive powers of the GB invariant. We will also assume that scalarization happens in a finite window of BH masses, centered at supermassive scales ($\sim 10^{9}\,M_{\odot}$), in light of the discussion above.
@@ II.2 Scalar charge model
In order to provide a description of the signal emitted by a population of SMBHBs, we need to model the dependence of the scalar charge on the BH mass. This procedure is possible as in EsGB we are dealing with secondary BH hairs. In other words, the scalar charge is not an independent parameter, but rather a function of mass $M$ and scalar field at infinity $\varphi_{\infty}$. Of course, this dependence is strongly model-specific, is likely to be affected by the BH spin and, in general, has no closed analytic form. We then follow a phenomenological approach, motivated by the discussion in the previous section. In more detail, we model the scalar charges with the functional form
@@
where $\Theta_{\text{H}}$ is the Heaviside step-function and we defined for later convenience $M_{9}\equiv M/(10^{9}M_{\odot})$ and $\lambda_{9}\equiv\lambda/(10^{9}M_{\odot})$. This curve is represented in Fig. 1. Notice that the model displays an upper mass bound at the point at which the function $\tilde{\alpha}$ becomes negative. As far as we now, such upper mass bound is present in all scalarization models in the literature. On the other hand, the Heaviside function produces a sharper cutoff at lower masses. In our analysis, we adopt the value of the cutoff in Eq. (10) as a proxy for the model-dependent scalarization threshold identified in Ref. [18]. This choice is intentionally made to emphasize the potential impact of scalarization on the SGWB by allowing for a wider scalarization window, while preserving the qualitative features of the underlying theory. Choosing higher cutoff values, consistent with Fig. 2 of Ref. [18], would yield a narrower window and a correspondingly weaker imprint on the PTA signal. To highlight the observational signatures, we focus on cutoff values that yield detectable effects in current PTA data. Future observations, with longer time spans and improved precision, will be able to test more restrictive thresholds. Finally, notice that care should be taken when comparing Fig. 1 to the results in Ref. [18], as we are representing the dimensionless scalar charge, rather then the dimensionful scalar charge divided by $\lambda$.
@@ III Spectral shape of the SGWB in EsGB
[…]
@@
$\delta t_{a}(t)\delta t_{b}(t)\right>\sim\xi(\theta_{ab})\int_{f_{l}}^{f_{h}}{\rm d}f\,P(f),$ (11)
@@ III.1 Comoving merger density
In order to compute the characteristic strain in Eq. (14), we need to specify the comoving merger density, which informs us about the underlying astrophysical population of SMBHBs. This quantity is among the main outputs of semi-analytic models, see e.g. [61, 62, 63, 64, 65]. In this work, we adapt the analytic merger density prescription developed in Ref. [66], which successfully captures the key features of semi-analytic models, while remaining entirely agnostic about the framework-specific merger history of SMBHBs. This allows us to focus exclusively on the aspects directly constrained by PTA data, and not those that require assumptions about a specific merger history. Notice that semi-analytic models usually track the evolution of dark matter merger trees, and then empirically relate the merger density of galactic halos to the merger density of SMBHBs. Thus, as long as the scalar coupling to matter is negligible, the astrophysical model developed in Ref. [66] is totally insensitive to the underlying theory of gravity.
@@
with $H_{0}$ the Hubble constant, $\Omega_{m,0}$ and $\Omega_{\Lambda,0}$ the matter and dark energy densities at present times. As a benchmark, we take $H_{0}=70\,\text{km s}^{-1}\text{Mpc}^{-1}$, $\Omega_{m,0}=0.3$ and $\Omega_{\Lambda,0}=0.7$. We can factor Eq. (15) in an overall amplitude, with $\text{Mpc}^{-3}$ dimensions, a redshift-dependent term and a mass-dependent term. Moreover, we consider a normalized distribution $p(\log_{10}q)$, to account for the fact that different systems may have different mass ratios $q$. In particular, we take it to be a log-uniform distribution in the range $\log_{10}q\in[-1.5,0]$, which is broadly in agreement with previous works [67, 68, 69]. As we will show later in Fig. 6, using a uniform distribution in the same mass ratio range, namely $q\in[0.03,1]$, does not appreciably impact the results. However, notice that if the mass ratio distribution is strongly peaked around $q=1$, the effect of EsGB gravity will be significantly suppressed. In this limit, for nearly equal-mass binaries, the difference in the scalar charges of the two components becomes negligible, leading to a strong suppression of the dipolar scalar emission (see Eq. (49) below). In the following, we will proceed under the caveat that the true astrophysical distribution does not strongly favor equal mass binaries. Therefore, unless otherwise specified, we will just assume $\log_{10}q\in[-1.5,0]$, as described before. We now sum up the priors on the parameters of Eq. (15), with a short description of each of them:

### v2

@@ Abstract
Scalar-tensor theories with a scalar field coupled to the Gauss-Bonnet invariant can evade no-hair theorems and allow for non-trivial scalar profiles around black holes. This coupling is characterized by a length scale $\lambda$, which, in an effective field theory perspective, sets the threshold below which deviations from General Relativity become significant. LIGO/VIRGO constraints indicate $\lambda$ is small, implying supermassive black holes should not scalarize. However, recent work suggests that scalarization can occur within a narrow window of masses, allowing supermassive black holes to scalarize, while leaving LIGO/VIRGO sources unaffected. We explore the impact of this scenario on the stochastic gravitational wave background recently observed by Pulsar Timing Arrays. We find that scalarization can alter the characteristic strain produced by circularly inspiralling supermassive black hole binaries and that current data shows a marginal preference for a non-zero $\lambda$. However, similar signatures could arise from astrophysical effects such as orbital eccentricity or environmental interactions, emphasizing the need for improved modeling and longer observations to discriminate among the different scenarios.
@@ I Introduction
On the other hand, it has been shown that most of the interaction terms admitted in (ghost-free) ST gravity do not affect black hole (BH) solutions at the background level. In other words, BHs in most ST theories are exactly the same as in GR, meaning that they are fully determined by mass, spin and, in principle, electric charge, while the additional scalar DOF provides no further “charge” to the BH. This fact is related to the well-known no-hair conjecture and has been rigorously formalized in some specific cases through no-hair theorems [9, 10, 11, 12, 13].
Nevertheless, such theorems rely on a set of crucial assumptions that cannot always be taken for granted. In particular, it has been shown that a specific interaction term, among all those leading to second-order equations of motion (EOM), can violate all known no-hair theorems, allowing for BH solutions with a non-trivial scalar profile [14, 15]. This term is given by $\lambda^{2}F(\varphi)\mathcal{G}$, where $\lambda$ is a coupling constant with dimension of a length, $\varphi$ is the scalar field, $F$ is an analytic function, and $\mathcal{G}$ is the Gauss-Bonnet (GB) invariant
@@
The interesting properties of this model include BH solutions coupled to the scalar field via the so-called scalar charge [14, 15, 16, 17, 18], which has been used to constrain the coupling parameter $\lambda$ using Gravitational Waves (GWs) observations from the LIGO/Virgo collaboration [2].
@@
An interesting feature of this model is that the best constraints on $\lambda$ actually come from the smaller BHs from which we observe the GWs. This is tied to the dimensionful nature of the coupling: the smaller the BH scale, the better the constraint on $\lambda$. In fact, one can view interaction terms coming with a dimensionful coupling constant $\lambda$, such as the scalar-GB coupling, as irrelevant operators in an effective field theory (EFT) perspective [19]. This implies that the contribution of this operator is suppressed at scales that are much larger than $\lambda$, whereas deviations from GR are expected to become increasingly significant at smaller scales. For this reason, it is usually assumed that supermassive BHs (SMBHs) are the same as in GR, even in presence of the scalar-GB coupling.
@@
The EFT interpretation of ST theories can also be made more concrete, as it can be shown that some specific operators arise as the low-energy limit of quantum gravity theories. This is the case for some scalar-GB couplings, that emerge in the low-energy limit of string theories [20]. However, in principle, one can suppose that high-energy effects, at a given scale, lead BHs back to the GR configuration, providing a well-delimited range of scales in which deviations from GR can occur. This happens, for instance, in a recently proposed model [18], in which SMBHs can possess a scalar charge while lower mass BHs are the same as in GR. It is therefore reasonable to envision departures from GR in the signals that we receive from SMBHs, which, when coupled in binary systems, are considered the most likely source of the recently observed stochastic GW background (SGWB).
@@
Indeed, in recent years, Pulsar Timing Array (PTA) collaborations reported evidence for a temporally-correlated stochastic process leaving its imprint on the times of arrival (TOAs) of radio beams coming from a network of monitored millisecond pulsars in our Galaxy [21, 22, 23, 24, 25]. Furthermore, the correlations observed among different pulsars were found to be consistent with the Hellings-Downs (HD) pattern [26], as expected for a SGWB. Although the observed data can also be explained with Early Universe processes (see e.g. Refs. [27, 28]), the most likely interpretation consists in the incoherent superposition of GWs emitted by a population of inspiralling SMBH binaries (SMBHBs).
Assuming circular orbits — which we restrict to in the following — the characteristic strain spectrum associated with GW-driven orbital decay is expected to follow a power-law scaling as $f^{-n}$, with the spectral index $n=2/3$, where $f$ denotes the emitted GW frequency [29, 30]. Eccentric orbits as well as the presence of any additional non-GW energy loss channel - such as dynamical friction or, in our case, scalar radiation- would accordingly modify the spectral index $n$ [31, 32, 33, 34, 35, 36, 37]. On the other hand, the predicted angular correlation pattern is a property of the polarization of GWs at the Earth location, and will essentially coincide with the HD curve whenever - as we consider in the following - the scalar field is not coupled to matter.
@@
In this article, we hence study the modification of the SGWB predicted by an assembly of circular SMBHBs evolving under the ST theory of gravity studied in Ref. [18]. In Sec. II, we summarize the framework of black hole scalarization and lay the foundations for the analysis developed throughout the paper. In Sec. III, we review how to construct the SGWB from an astrophysical population of SMBHBs, and we compare the resulting characteristic strain in GR with that predicted by the ST theory described in Sec. II. In Sec. IV, we analyze the EPTA DR2New dataset, comprising the most recent 10.3 yr of data of the EPTA collaboration, and the NANOGrav 15yr data, looking for signatures of non trivial SMBHBs scalarization. Finally, in Sec. V, we summarize our findings and conclude with a brief discussion of their implications and possible directions for future work. Additional technical details and supplementary plots are provided in the Appendices. In the following, we will work in natural units $c=1$, while keeping expressed the gravitational coupling constant $G$, and using the mostly plus signature of the metric tensor (,+,+,+)(-,+,+,+).
@@ II.1 Motivation
Notice that the shift-symmetric models exhibit characteristic features, like finite-radius singularities, whose location depends on the coupling $\lambda$ [14, 15]. The presence of these singularities, if one requires them to be always located within the event-horizon, imposes a lower bound on the BH size, that translates into an observational constraint on the coupling $\lambda$ for this theory [42, 43]. However, it is worth mentioning that the shift-symmetric EsGB theory, when requiring that it produces observable deviations from GR at astrophysical scales, presents issues with causality [44]. In the dilatonic case, one also encounters a lower BH mass limit, set by the dimensionful coupling $\lambda$, so that the best experimental constraints can be put from low mass BH observations [45].
@@
with $\tilde{\lambda}$ being another dimensionful coupling constant. In this model, to ensure that the Schwarzschild BH still solve the EOMs, one has to restrict to the case $\varphi=0$. In the $\tilde{\lambda}=0$ case, the Schwarzschild solution is admitted with any constant $\varphi_{0}$, as $\mathcal{G}$ yields a trivial contribution to the EOM, being it a surface term. In the more general case $\tilde{\lambda}\neq 0$, the condition for the tachyonic instability to occur is now given by
@@
This practically demonstrates that higher-order terms in EsGB can, in principle, protect BHs from scalarization at small scales, where the standard exponential model described in Ref. [16] is affected by hyperbolicity loss, while allowing for this phenomenon to occur at larger scales. For this reason, the constraints on scalar charges provided by LIGO-Virgo do not necessarily imply that scalarization should be excluded for SMBHs.
In the following, we will be agnostic about the specific model,  [^3] Notice that, in a work released after our initial submission, the model presented in Ref. [18] has been argued to be inconsistent from an EFT point of view [49]. We stress that our reasoning remains valid regardless of the specific model, as long as scalarization takes place at supermassive scales. and just assume that the scalar field is coupled to some positive powers of the GB invariant. We will also assume that scalarization happens in a finite window of BH masses, centered at supermassive scales ($\sim 10^{9}\,M_{\odot}$), in light of the discussion above.
@@ II.2 Scalar charge model
In order to provide a description of the signal emitted by a population of SMBHBs, we need to model the dependence of the scalar charge on the BH mass. This procedure is possible as in EsGB we are dealing with secondary BH hairs. In other words, the scalar charge is not an independent parameter, but rather a function of mass $M$ and scalar field at infinity $\varphi_{\infty}$. Of course, this dependence is strongly model-specific, is likely to be affected by the BH spin and, in general, has no closed analytic form. We then follow a phenomenological approach, motivated by the discussion in the previous section.
In more detail, we model the scalar charges with the functional form
@@
where $\Theta_{\text{H}}$ is the Heaviside step-function and we defined for later convenience $M_{9}\equiv M/(10^{9}M_{\odot})$ and $\lambda_{9}\equiv\lambda/(10^{9}M_{\odot})$. This curve is represented in Fig. 1. Notice that the model displays an upper mass bound at the point at which the function $\tilde{\alpha}$ becomes negative. As far as we now, such upper mass bound is present in all scalarization models in the literature. On the other hand, the Heaviside function produces a sharper cutoff at lower masses.
Our choice of the functional form in Eq. (10) is intended as a minimal model that captures the key features of scalarization at the SMBH scale discussed so far: namely, the existence of a mass window bounded both from below and above, along with a maximum scalar charge at intermediate scales. We emphasize that the adopted parametrization is not meant to reproduce the precise dynamics of any specific ST theory, but rather to serve as a flexible template for studying potential observational imprints in the the PTA data, as we will do below.
Moreover, in our analysis, we adopt the value of the cutoff in Eq. (10) as a proxy for the model-dependent scalarization threshold identified in Ref. [18]. This choice is intentionally made to emphasize the potential impact of scalarization on the SGWB by allowing for a wider scalarization window, while preserving the qualitative features of the underlying theory. Choosing higher cutoff values, consistent with Fig. 2 of Ref. [18], would yield a narrower window and a correspondingly weaker imprint on the PTA signal. To highlight the observational signatures, we focus on cutoff values that yield detectable effects in current PTA data. Future observations, with longer time spans and improved precision, will be able to test more restrictive thresholds. Finally, notice that care should be taken when comparing Fig. 1 to the results in Ref. [18], as we are representing the dimensionless scalar charge, rather then the dimensionful scalar charge divided by $\lambda$.
@@ III Spectral shape of the SGWB in EsGB
PTAs are facilities timing an ensemble of Galactic pulsars for decades, with an approximate cadence of $\sim 1\,\text{week}$. The inverse of the observational timespan sets the minimum frequency PTAs are sensitive to, roughly $1\,\text{nHz}$, while the inverse cadence sets the maximum frequency, around $1\,\mu\text{Hz}$.
@@
$\left<\delta t_{a}(t)\delta t_{b}(t)\right>\sim\xi(\theta_{ab})\int_{f_{l}}^{f_{h}}{\rm d}f\,P(f),$ (11)
@@ III.1 Comoving merger density
In order to compute the characteristic strain in Eq. (14), we need to specify the comoving merger density, which informs us about the underlying astrophysical population of SMBHBs. This quantity is among the main outputs of semi-analytic models, see e.g. [62, 63, 64, 65, 66].
In this work, we adapt the analytic merger density prescription developed in Ref. [67], which successfully captures the key features of semi-analytic models, while remaining entirely agnostic about the framework-specific merger history of SMBHBs. This allows us to focus exclusively on the aspects directly constrained by PTA data, and not those that require assumptions about a specific merger history. Notice that semi-analytic models usually track the evolution of dark matter merger trees, and then empirically relate the merger density of galactic halos to the merger density of SMBHBs. Thus, as long as the scalar coupling to matter is negligible, the astrophysical model developed in Ref. [67] is totally insensitive to the underlying theory of gravity.
@@
with $H_{0}$ the Hubble constant, $\Omega_{m,0}$ and $\Omega_{\Lambda,0}$ the matter and dark energy densities at present times. As a benchmark, we take $H_{0}=70\,\text{km s}^{-1}\text{Mpc}^{-1}$, $\Omega_{m,0}=0.3$ and $\Omega_{\Lambda,0}=0.7$. We can factor Eq. (15) in an overall amplitude, with $\text{Mpc}^{-3}$ dimensions, a redshift-dependent term and a mass-dependent term. Moreover, we consider a normalized distribution $p(\log_{10}q)$, to account for the fact that different systems may have different mass ratios $q$. In particular, we take it to be a log-uniform distribution in the range $\log_{10}q\in[-1.5,0]$, which is broadly in agreement with previous works [68, 69, 70]. As we will show later in Fig. 6, using a uniform distribution in the same mass ratio range, namely $q\in[0.03,1]$, does not appreciably impact the results. However, notice that if the mass ratio distribution is strongly peaked around $q=1$, the effect of EsGB gravity will be significantly suppressed. In this limit, for nearly equal-mass binaries, the difference in the scalar charges of the two components becomes negligible, leading to a strong suppression of the dipolar scalar emission (see Eq. (51) below). In the following, we will proceed under the caveat that the true astrophysical distribution does not strongly favor equal mass binaries. Therefore, unless otherwise specified, we will just assume $\log_{10}q\in[-1.5,0]$, as described before.
We now sum up the priors on the parameters of Eq. (15), with a short description of each of them: